# When Large Language Models Meet UAVs: How Far Are We?

### v1

@@ 4.1. Data Collection and Processing
In the data collection process, we surveyed GitHub to identify industry projects integrating LLMs with UAVs. Specifically, we conducted keyword-based retrieval using two curated sets of keywords covering common terms related to UAVs (e.g., “drone”, “UAV”, “quadcopter”) and LLMs (e.g., “GPT”, “large language model”, “LLM”). We retained repositories where at least one UAV-related keyword and one LLM-related keyword co-occurred in the title, README, or description. Subsequently, two authors independently reviewed each candidate repository to verify its relevance. Ultimately, we identified 56 projects that provided either complete engineering solutions or prototype demonstrations of UAV-LLM integration.
@@
To investigate the gap between academia and industry in the integration of LLMs with UAVs, we conducted a comparative analysis based on the task classification framework of LLM applications in UAVs established in RQ1. The academic data source consists of 74 relevant publications obtained through a systematic literature review (SLR) conducted in RQ1, while the industry data source comprises 56 real-world projects collected by two researchers from major technical and social platforms. The task classification framework was developed by the two researchers in RQ1 using an open card sorting method, which distilled nine types of tasks performed by LLMs in UAV scenarios and five core application domains. We annotated and categorized each data point in both datasets according to the tasks and application directions of LLMs, mapping them to the four core domains and nine specific tasks shown in Figure 1. The frequency distribution of each category within the academic and industrial ecosystems was then analyzed. This entire process involved multiple rounds of iteration and discussion between the two researchers to reach consensus, ensuring scientific rigor and consistency in the results. The outcome provides a solid foundation for identifying potential commonalities and differences between academia and industry in this emerging field.
Figure 3. Statistics on the application of LLMs in UAVs in academia and industry.
4.2. Results
4.2.1. Current State of Industry
@@
Firstly, we conducted a statistical analysis of the tasks performed by LLMs in UAV systems for each industrial project. The results include: Natural Language Command Parsing (29/56, 51.79%), Multimodal Information Parsing (17/56, 30.36%), single-UAV Task Reasoning and Planning (28/56, 50%), Multi-UAV Task Reasoning and Planning (8/56, 14.29%), RAG-based Decision Support (5/56, 8.92%), Control Logic and Command Generation (40/56, 71.43%), Control Logic and Command Optimization (2/56, 3.57%), Information Summarization and Report Generation (12/56, 21.43%), and LLM-based Chatbot (8/56, 14.29%). We observed that developers are most focused on flight control in UAV systems, with projects performing the task of Control Logic and Command Generation accounting for 71.43% of all projects. Natural Language Command Parsing and Single-UAV Task Reasoning and Planning are also popular, accounting for 51.79% and 50%, respectively, indicating an increasing demand in the industry for task automation and intelligent interaction.
Finding 3: In over 50% of the projects, the tasks handled by large models are concentrated in areas such as flight control, task planning, and human-machine interaction, with less than 10% of the projects focusing on other tasks.
@@
In contrast, tasks such as RAG-based Decision Support and Control Logic and Command Optimization are less commonly applied, which could be because these tasks require more external knowledge bases and datasets. Additionally, Multimodal Information Parsing and Multi-UAV Task Reasoning and Planning have seen some applications in the industry but still face challenges related to real-time performance, computational resource demands, and the inherent complexity of these tasks.
4.2.2. The differences between academia and industry
@@
We analyzed the literature and projects that integrate UAVs with LLMs, presenting their specifics in Figure 3 and identifying three key differences between academia and industry in the integration of UAVs with LLMs. In Figure. 3, “UAV Workflow” and “Tasks performed by LLMs in UAVs” represent the four UAV workflows and the nine tasks performed by LLMs in UAV systems that we identified in RQ1. “(Academia) Literature ID” and “(Industry) Project ID” refer to the IDs of the literature and projects we collected. We have analyzed the UAV tasks involved in each of the literature and projects.
@@
Prevalence and Distribution of LLM-UAV projects. In terms of prevalence, there are currently fewer industrial projects combining large language models with UAVs than academic research, which could be attributed to the early development stage and high development cost of LLM-UAV integration. The focus range of academic research on the nine tasks spans from 12.16% to 39.19%, while in industry, the range is from 3.57% to 71.43%. In terms of scope and depth, this difference indicates that academic research has a more dispersed focus across various tasks, with a broad research scope and no apparent concentration of high attention on specific tasks. Instead, it covers multiple key areas such as multi-UAV collaboration, decision support, and environmental perception. In contrast, industrial projects exhibit a trend of task centralization, with higher attention given to tasks directly related to product interaction and execution, focusing more on practicality and the automation of task execution.
@@
Application Fields. Academic research focuses on cutting-edge areas, emphasizing theoretical modeling and task optimization, including algorithm optimization, data fusion, and the application of external knowledge. For example, academic research addresses multi-UAV collaboration and decision support using RAG and reinforcement learning techniques (Liu et al., 2024; Javaid et al., 2024), combines Q-Learning, graph neural networks, and LLM reasoning to achieve dual optimization of communication and task execution (Li et al., 2025b), and utilizes LLMs for link quality prediction and dynamic resource allocation (Tang et al., 2024). On the other hand, the industrial sector faces higher real-time requirements and emphasizes practical operability and scalability, focusing on how to improve task execution efficiency and control accuracy in real-world applications. For example, in more than 50% of the projects, large models are responsible for tasks such as Natural Language Command Parsing, Control Logic and Command Generation, and Single-UAV Task Reasoning and Planning. These tasks are concentrated in areas like flight control, task planning, and human-machine interaction.
Finding 4: Regarding the development of RAG-based Decision Support and Control Logic and Command Optimization, academic research is primarily focused on these areas, with respective proportions of 20.27% and 12.16%. In contrast, their application in industrial projects is relatively limited, with respective proportions of 8.92% and 3.57%.
@@
Different Integration Modes Between LLMs and UAVs. To gain a deeper understanding of the technical details of LLMs’ application in the UAV field in both academia and industry, we identified three main integration modes of LLMs with UAVs based on the location of the LLM and the interaction between the LLM and UAV. These modes are: External (Offboard Invocation) mode, Hybrid (Onboard Assistance) mode, and Internal (Onboard Control) mode. We researched their distribution in current academic studies and industrial projects.
@@
    Offboard Invocation (External): In this mode, the LLM is deployed outside the UAV, either in the cloud or at a ground-based edge computing node. It receives task information from the UAV via remote APIs and returns task suggestions and semantic interpretations. This mode offers flexible deployment and strong computational power but suffers from high communication latency and limited real-time performance, making it more suitable for non-real-time tasks such as large-scale inference and complex semantic parsing. The use of this mode in academia (52.7%) is higher than in industry (41.07%), reflecting that academic research is still largely focused on cloud-based invocation and experimental validation, with tasks concentrated on algorithm verification and offline task planning. In contrast, industry has reduced the proportion of offboard invocation, indicating that some tasks have migrated to edge/onboard systems to reduce latency and improve real-time performance.
@@
    Onboard Assistance (Hybrid): In this mode, the LLM is deployed on an edge computing device within the UAV (such as Jetson, RK3588, etc.). It does not directly generate control commands but acts as a decision-support module, collaborating with the UAV’s flight control system to support task planning, path inference, and multi-UAV coordination. This mode balances intelligence and stability, making it suitable for complex tasks such as real-time perception, path adjustment, and local decision-making. The usage of this mode in industry (50%) is higher than in academia (38%), suggesting that in real-world industrial deployments, real-time performance and stability are considered more important than extreme inference capabilities. Academic research, however, may be limited by hardware constraints and deployment complexity.
@@
    Onboard Control (Internal): In this mode, the LLM is deeply integrated into the UAV’s flight control system, directly responsible for task decomposition, action decision-making, and control command generation. Its advantage is that the UAV’s flight control directly relies on the LLM, providing good real-time performance. It also offers high real-time capability and autonomy, making it a key architecture for achieving fully autonomous flight. However, the disadvantage is the high demand for local computational resources, which is limited by the model size. The usage of this mode is low in both academia (5.4%) and industry (10.71%), reflecting a cautious attitude. This mode is currently more commonly used in high-value, specific tasks with clear ROI (such as rescue, defense, and mining). It can be speculated that, as technology advances, particularly with edge computing and lightweight models, this mode will have significant growth potential in the future.
Finding 5: In the three integration methods of LLMs and UAVs, academia most commonly adopts External-offboard invocation(52.7%), while industry tends to prioritize Hybrid-onboard assistance (50%). This indicates that academia focuses more on algorithm validation and performance evaluation in controlled environments, whereas industry places greater emphasis on real-time performance and the practical application of resource constraints. .
Finding 6: Currently, LLMs are rarely directly involved in UAV control, but the development of edge computing and lightweight models may bring new possibilities for the application of large models in UAVs.
@@
5. RQ3: How do real-world practitioners perceive these differences, and what factors contribute to them?
@@
Section 4 presents preliminary findings that reveal the differences in UAV plus LLM integration between academia and industry. To further explore the potential reasons behind these differences and identify possible solutions to bridge the gap, we conducted an online survey targeting real-world UAV developers. The goal was to gain deeper insights into practitioners’ perspectives on integrating LLMs into UAV systems, as well as to examine the key obstacles and design choices encountered in this integration process.
5.1. Survey Design
The UAV survey is divided into four sections.
@@ 5.1. Survey Design
Demographics. (Q1-Q4) This section collected participants’ professional background to ensure relevance and interpretability of their responses. Participants were asked about their primary domain (e.g., UAVs, robotics, LLMs) (Q1), responsibility (hardware/software/algorithm development, testing, research, etc.) (Q2), years of experience (Q3), and geographical location (Q4).
@@
Practices and Challenges. (Q5-Q6) We explored whether the practitioners had ongoing projects combining LLMs and UAVs (Q5), and if not, what challenges (e.g., cost, technical limitations, regulation, unclear business models) had prevented integration (Q6).
@@
Application Maturity and Integration Approach. (Q7-Q10) Participants were asked to rate the maturity of various LLM application scenarios in UAVs, such as Natural Language Command Parsing, Multimodal Information Parsing, Single-UAV Task Reasoning and Planning, and Multi-UAV Task Reasoning and Planning, as well as other related tasks (Q7). Additional open-ended questions inquired about other potential applications desired by practitioners (Q8), the implementation approaches of UAV-LLM integration(e.g., External, Hybrid, and Internal) (Q9), and the LLM’s capabilities desired for UAV tasks (Q10).
@@
Industry-Academia Comparison. (Q11-Q13) To understand the alignment between research and practice, we asked whether participants referred to academic work in their projects, and collected their views on the differences between academic and industrial efforts (Q11). Specific factors such as development lifecycles, technical requirements (Q12), and integration strategies were covered. Furthermore, we investigated the perceived underlying reasons for these differences, such as real-world requirements for technical maturity, performance, safety, and cost (Q13).
5.1.1. Participant Recruitment.
@@
Our target practitioners for the survey were UAV developers. Our goal was to distribute the survey to GitHub developers who had contributed to UAV-related projects. First, we searched for UAV-related repositories on GitHub using keywords such as “UAV”. Then, we used the GitHub Search API to collect the email addresses and names of contributors to the selected repositories. In total, we obtained 802 developer email addresses and sent out invitations to participate in the survey. Additionally, we shared the survey with several industry collaborators working at well-known companies, such as China Mobile Communications Corporation, with a total of six industry partners involved. Since some developers might not be familiar with the term “UAV,” we provided a brief explanation of the concept in the email to avoid confusion.
5.1.2. Data Analysis.
@@
We received a total of 55 valid responses from 15 different countries, with a response rate of approximately 6.85%. Among these practitioners, we excluded three individuals who indicated that they had no development experience related to UAVs, robotics, or LLMs, and conducted further analysis on the remaining 52 participants. The average experience of the remaining practitioners was 5.9 years. Their primary responsibilities included hardware development (12, 23.01%), software development (37, 71.15%), algorithm development (25, 48.08%), testing (9, 17.31%), and research (13, 25%). Note that a practitioner may concurrently hold multiple roles, such as developing and testing software.
5.2. Results
5.2.1. Practices and Challenges in UAV-LLM integration
@@
Real-world Practices. In the survey, 40.4% of the practitioners have started to advance projects that integrate LLMs with UAVs), with 19.05% of them having fully deployed the technology, and 80.95% of them being in the proof-of-concept phase. This indicates that, although the integration of LLMs with UAVs is still in the exploratory stage, some practitioners have already successfully implemented the initial application of this technology. However, 59.6% of the practitioners have not yet initiated this project. For those practitioners who have not implemented the project, we further inquired about the specific reasons for not advancing this technology. The survey revealed the main challenges.
@@
Challenges Perceived by Practitioners. The analysis shows that the obstacles mentioned by practitioners are mainly related to technology, resources, and market factors. In particular, regarding technology, 54.8% of the practitioners believe that existing technologies already meet their needs, and therefore, they are not in a rush to introduce new technological solutions. This suggests that in some traditional UAV application scenarios, the existing technology and control systems are mature enough to reliably complete tasks without the need for more complex LLM technology. 51.6% of practitioners believe that existing large models cannot meet performance requirements, reflecting the current shortcomings of LLMs in terms of performance, computational power, and real-time capabilities, which are still insufficient to meet the high-performance demands in the UAV field. Additionally, 45.2% of practitioners pointed out the security risks of large models, with some noting that “LLMs are not being used on board in safety-critical applications” and “LLMs should not be on a critical system due to their ‘black box’ nature.” This indicates that data privacy, as well as the transparency and interpretability of models, remain major obstacles to the widespread application of LLMs in the UAV domain.
@@
In addition to technological issues, 25.8% of practitioners mentioned the high costs, noting that the high computational resource demands and related hardware limitations make it difficult for these practitioners to accept when considering cost-effectiveness. 22.6% of practitioners stated that the lack of training data or fine-tuning methods limits their application, preventing them from effectively utilizing the existing large models. Furthermore, 29.0% of practitioners believe that the lack of relevant knowledge and technology is another reason for not advancing this project.
Finding 7: Currently, 40.4% of practitioners have attempted to integrate LLMs with UAVs in their projects. The main reasons for not attempting are “Existing technology has met the project requirements” (54.8%), “Existing large language models cannot meet performance requirements” (51.6%), and “Security risks of large language models” (45.2%).
Finding 8: Some practitioners also expressed doubts about the application of LLMs in critical systems such as UAVs. This suggests that the integration of LLMs with UAVs is still in the early stages of exploration. Although some practitioners have already started applying them, there is no urgent demand from the industry, and they face significant technical obstacles. To advance this field, it is necessary to enhance the performance, scalability, and reliability of LLMs in critical tasks.
Table 2. Practitioners’ perceptions on tasks maturity
@@
Application Maturity Perceived by Practitioners. To assess the current industrial acceptance and technical maturity of large models in UAV tasks, we invited practitioners to rate the technical maturity of various tasks of LLMs in UAVs. We used the Likert scale with a scale from 1 to 5, with 1 indicating “completely immature,” 3 representing neutral, and 5 indicating “very mature”. The results, as shown in  2, indicate that the average score for the technical maturity of all tasks is below 3, indicating that while LLMs hold significant application potential in these areas, they also face numerous challenges. Among these, practitioners considered LLMs to have relatively higher technical maturity in Natural Language Command Parsing and Information Summarization and Report Generation, with scores of 2.89 and 2.96, respectively. These tasks involve basic processing of language and data, and existing LLMs can provide effective support in these areas, resulting in a relatively higher technical maturity. However, for more complex tasks, such as Multi-UAV Task Reasoning and Planning (score 2.40) and Control Logic and Command Generation (score 2.51), the technical maturity was lower. This suggests that, while multi-UAV coordination and autonomous control theoretically have significant potential, there are still many technical obstacles in effectively applying LLMs to these complex tasks in practical industrial applications. These challenges need further improvement in areas such as accuracy, real-time performance, and collaborative computing capabilities.
@@
Additionally, two practitioners specifically highlighted potential application scenarios for LLMs in the UAV field, including: supporting users and new developers in setting up UAV systems (e.g., helping configure ArduPilot Wiki) and applications in areas such as natural disaster response, maritime accidents, and search and rescue (SAR), environmental monitoring and hazard detection, and urban safety and public monitoring. This highlights the potential for LLMs in the UAV field in practical industrial applications, particularly in high-complexity and multi-task processing scenarios, which remain to be further explored.
Finding 9: In terms of technology maturity perceived by practitioners, Natural Language Command Parsing (score 2.89) and Information Summarization and Report Generation (2.96 points) received relatively higher scores, while Multi-UAV Task Reasoning and Planning (score 2.40) and Control Logic and Command Generation (score 2.51) scored lower. This indicates that current LLMs can provide effective support for tasks involving basic language and data processing, but tasks that require higher accuracy, real-time performance, and computational resources still need further improvement to meet the decision support and real-time response requirements of existing UAVs.
@@
Common Integration Approaches. To gain a deeper understanding of the technical details of integrating LLMs with UAVs, we explored the main integration methods of LLMs in the UAV field. Specifically, we asked practitioners to select from various integration methods. The survey results show that Hybrid-Onboard Assistance is the most common method, with 82.7% of practitioners choosing this option. In this approach, LLMs typically do not directly control the UAV’s flight but serve as a supporting tool for task planning, path reasoning, and decision support. This suggests that, currently, LLMs in the UAV field are more focused on enhancing decision-making capabilities to improve the intelligence of task execution, rather than directly intervening in the flight control system.
@@
In comparison, External-Offboard Invocation (i.e., deploying large models on cloud or ground devices to provide decision support via external interfaces) and Internal-Onboard Control (i.e., embedding large models into the UAV’s control system to generate flight control commands directly) were selected by 67.3% and 26.9% of practitioners, respectively. The application of these two methods is relatively limited and, in practical use, is not as common as the Hybrid-Onboard Assistance approach. This reflects that the application of LLMs in the UAV field is still in the early stages, with the industry more reliant on enhancing decision support rather than direct control.

### v2

@@ 4.1 Data Collection and Processing
In the data collection process, we surveyed GitHub to identify real-world projects integrating LLMs with UAVs. Specifically, we conducted keyword-based retrieval using two curated sets of keywords covering common terms related to UAVs (namely “drone”, “UAV”, “quadcopter”, “autopilot”, “PX4”, “multirotor”) and LLMs (namely “GPT”, “large language model”, “LLM”, “LLaMA”, “ChatGPT”, “transformer”). We retained repositories where at least one UAV-related keyword and one LLM-related keyword co-occurred in the title, README, or description. Subsequently, two authors independently reviewed each candidate repository to verify its relevance. Ultimately, we identified 56 projects that provided either complete engineering solutions or prototype demonstrations of UAV-LLM integration.
@@
To investigate the gap between researchers’ and developers’ practices in the integration of LLMs with UAV projects, we conducted a comparative analysis based on the task classification framework of LLM applications in UAVs established in RQ1. The dataset on researchers’ practices consists of 74 relevant publications obtained through a systematic literature review (SLR) conducted in RQ1, while the dataset on developers’ practices comprises 56 real-world projects collected by two researchers from major technical and social platforms. The task classification framework was developed by the two researchers in RQ1 using an open card sorting method, which distilled nine types of tasks performed by LLMs in UAV scenarios and five core application domains. We annotated and categorized each data point in both datasets according to the tasks and application directions of LLMs, mapping them to the four core domains and nine specific tasks shown in Figure 1. The frequency distribution of each category within the researcher and developer ecosystems was then analyzed. This entire process involved multiple rounds of iteration and discussion between the two researchers to reach consensus, ensuring scientific rigor and consistency in the results. The outcome provides a solid foundation for identifying potential commonalities and differences between researchers and developers in this emerging field.
Figure 3: Statistics on LLM applications in UAVs between researchers and developers.
4.2 Results
4.2.1 Current Practices
@@
Firstly, we conducted a statistical analysis of the tasks performed by LLMs in UAV projects for each developer. The results include: Natural Language Command Parsing (29/56, 51.79%), Multimodal Information Parsing (17/56, 30.36%), single-UAV Task Reasoning and Planning (28/56, 50%), Multi-UAV Task Reasoning and Planning (8/56, 14.29%), RAG-based Decision Support (5/56, 8.92%), Control Logic and Command Generation (40/56, 71.43%), Control Logic and Command Optimization (2/56, 3.57%), Information Summarization and Report Generation (12/56, 21.43%), and LLM-based Chatbot (8/56, 14.29%). We observed that developers are most focused on flight control in UAV projects, with projects performing the task of Control Logic and Command Generation accounting for 71.43% of all projects. Natural Language Command Parsing and Single-UAV Task Reasoning and Planning are also popular, accounting for 51.79% and 50%, respectively, indicating an increasing demand among developers for task automation and intelligent interaction.
[…]
@@
In contrast, tasks such as RAG-based Decision Support and Control Logic and Command Optimization are less commonly applied, which could be because these tasks require more external knowledge bases and datasets. Additionally, Multimodal Information Parsing and Multi-UAV Task Reasoning and Planning have seen some applications in the projects but still face challenges related to real-time performance, computational resource demands, and the inherent complexity of these tasks.
4.2.2 The differences between researchers and developers
@@
We analyzed the literature and projects that integrate UAVs with LLMs, presenting their specifics in Figure LABEL:fig:fig15 and identifying three key differences between researchers’ and developers’ practices in integrating UAV projects with LLMs. In Figure. LABEL:fig:fig15, “UAV Workflow” and “Tasks performed by LLMs in UAVs” represent the four UAV workflows and the nine tasks performed by LLMs in UAV projects that we identified in RQ1. “(Researchers) Literature ID” and “(Developers) Project ID” refer to the IDs of the literature and projects we collected. We have analyzed the UAV tasks involved in each of the literature and projects.
@@
Prevalence and Distribution of Real-World LLM-UAV projects. In terms of prevalence, there are currently fewer developers’ projects combining large language models with UAVs than academic research, which could be attributed to the early development stage and high development cost of LLM-UAV integration. The focus range of academic research on the nine tasks spans from 12.16% to 39.19%, while in projects, the range is from 3.57% to 71.43%. In terms of scope and depth, this difference indicates that researchers have a more dispersed focus across various tasks, with a broad research scope and no apparent concentration of high attention on specific tasks. Instead, it covers multiple key areas such as multi-UAV collaboration, decision support, and environmental perception. In contrast, developers’ projects exhibit a trend of task centralization, with higher attention given to tasks directly related to product interaction and execution, focusing more on practicality and the automation of task execution.
@@
Application Fields. Researchers focus on cutting-edge areas, emphasizing theoretical modeling and task optimization, including algorithm optimization, data fusion, and the application of external knowledge. For example, researchers address multi-UAV collaboration and decision support using RAG and reinforcement learning techniques [liu2024navagent, javaid2024large], combine Q-Learning, graph neural networks, and LLM reasoning to achieve dual optimization of communication and task execution [15], and utilize LLMs for link quality prediction and dynamic resource allocation [69]. On the other hand, developers’ projects face higher real-time requirements and emphasize practical operability and scalability, focusing on improving task execution efficiency and control accuracy in real-world applications. For example, in more than 50% of the projects, large models are responsible for tasks such as Natural Language Command Parsing, Control Logic and Command Generation, and Single-UAV Task Reasoning and Planning. These tasks are concentrated in areas like flight control, task planning, and human-machine interaction.
[…]
@@
Different LLM-UAV Integration Practices. To gain a deeper understanding of the technical details of how researchers and developers technically integrate LLMs into UAV projects, we identified three main integration modes of LLMs with UAVs based on the location of the LLM and the interaction between the LLM and UAV. These modes are: External (Offboard Invocation) mode, Hybrid (Onboard Assistance) mode, and Internal (Onboard Control) mode. We researched their distribution in current researchers’ and developers’ practices.
@@
    Offboard Invocation (External): In this mode, the LLM is deployed outside the UAV, either in the cloud or at a ground-based edge computing node. It receives task information from the UAV via remote APIs and returns task suggestions and semantic interpretations. This mode offers flexible deployment and strong computational power but suffers from high communication latency and limited real-time performance, making it more suitable for non-real-time tasks such as large-scale inference and complex semantic parsing. The use of this mode by researchers (52.7%) is higher than by developers (41.07%), reflecting that researchers are still largely focused on cloud-based invocation and experimental validation, with tasks concentrated on algorithm verification and offline task planning. In contrast, developers have reduced the proportion of offboard invocation, indicating that some tasks have migrated to edge/onboard systems to reduce latency and improve real-time performance.
@@
    Onboard Assistance (Hybrid): In this mode, the LLM is deployed on an edge computing device within the UAV (such as Jetson, RK3588, etc.). It does not directly generate control commands but acts as a decision-support module, collaborating with the UAV’s flight control system to support task planning, path inference, and multi-UAV coordination. This mode balances intelligence and stability, making it suitable for complex tasks such as real-time perception, path adjustment, and local decision-making. The usage of this mode by developers (50%) is higher than by researchers (38%), suggesting that in developers’ real-world projects, real-time performance and stability are considered more important than extreme inference capabilities. However, researchers are often limited by hardware constraints and deployment complexity. ß
@@
    Onboard Control (Internal): In this mode, the LLM is deeply integrated into the UAV’s flight control system, directly responsible for task decomposition, action decision-making, and control command generation. Its advantage is that the UAV’s flight control directly relies on the LLM, providing good real-time performance. It also offers high real-time capability and autonomy, making it a key architecture for achieving fully autonomous flight. However, the disadvantage is the high demand for local computational resources, which is limited by the model size. The usage of this mode is low among both researchers (5.4%) and developers (10.71%), reflecting a cautious attitude. This mode is currently more commonly used in high-value, specific tasks with clear ROI (such as rescue, defense, and mining). It can be speculated that, as capabilities improve, particularly with edge computing and lightweight models, this mode will have significant growth potential in the future.
@@
5 RQ3: How do real-world developers perceive these differences, and what factors contribute to them?
@@
Section 4 presents preliminary findings that reveal the differences in integrating UAVs with LLMs between researchers’ and developers’ practices. To further explore the potential reasons behind these differences and identify possible solutions to bridge the gap, we conducted an online survey targeting real-world UAV developers. The goal was to gain deeper insights into developers’ perspectives on integrating LLMs into UAV systems, as well as to examine the key obstacles and design choices encountered in this integration process.
5.1 Survey Design
The UAV survey, designed following the guidelines for personal opinion surveys proposed by Kitchenham and Pfleeger [kitchenham2008personal], adopts an anonymous format [tyagi1989effects] to encourage candid feedback and increase response rates. It is divided into four sections.
@@ 5.1 Survey Design
Demographics. (Q1-Q4) This section collected developers’ professional background to ensure relevance and interpretability of their responses. Developers were asked about their primary domain (e.g., UAVs, robotics, LLMs) (Q1), responsibility (hardware/software/algorithm development, testing, research, etc.) (Q2), years of experience (Q3), and geographical location (Q4).
@@
Practices and Challenges. (Q5-Q6) We explored whether the developers had ongoing projects combining LLMs and UAVs (Q5), and if not, what challenges (e.g., cost, technical limitations, regulation, unclear business models) had prevented integration (Q6).
@@
Application Maturity and Integration Approach. (Q7-Q10) Developers were asked to rate the maturity of various LLM application scenarios in UAVs, such as Natural Language Command Parsing, Multimodal Information Parsing, Single-UAV Task Reasoning and Planning, and Multi-UAV Task Reasoning and Planning, as well as other related tasks (Q7). Additional open-ended questions inquired about other potential applications desired by developers (Q8), the implementation approaches of UAV-LLM integration(e.g., External, Hybrid, and Internal) (Q9), and the LLM’s capabilities desired for UAV tasks (Q10).
@@
Academia-Industry Comparison. (Q11-Q13) To understand the alignment between research and practice, we asked whether developers referred to research work in their projects, and collected their views on the differences between academic and industrial efforts (Q11). Specific factors such as development lifecycles, technical requirements (Q12), and integration strategies were covered. Furthermore, we investigated the perceived underlying reasons for these differences, such as real-world requirements for technical maturity, performance, safety, and cost (Q13).
5.1.1 Developer Recruitment.
@@
Our target practitioners for the survey were UAV developers. Our goal was to distribute the survey to GitHub developers who had contributed to UAV-related projects. First, we searched for UAV-related repositories on GitHub using keywords such as “UAV”. Then, we used the GitHub Search API [github_api_docs] to collect the email addresses and names of contributors to the selected repositories. In total, we obtained 802 developer email addresses and sent out invitations to participate in the survey. Additionally, we shared the survey with several industry collaborators working at well-known companies, such as China Mobile Communications Corporation, with a total of six industry partners involved. Since some developers might not be familiar with the term “UAV,” we provided a brief explanation of the concept in the email to avoid confusion.
5.1.2 Data Analysis.
@@
We received a total of 55 valid responses from 15 different countries, with a response rate of approximately 6.85%. Among these developers, we excluded three individuals who indicated that they had no development experience related to UAVs, robotics, or LLMs, and conducted further analysis on the remaining 52 developers. The average experience of the remaining developers was 5.9 years. Their primary responsibilities included hardware development (12, 23.01%), software development (37, 71.15%), algorithm development (25, 48.08%), testing (9, 17.31%), and research (13, 25%). Note that a developer may concurrently hold multiple roles, such as developing and testing software.
5.2 Results
5.2.1 Practices and Challenges in UAV-LLM integration
@@
Real-world Practices. In the survey, 40.4% of developers have started advancing projects that integrate LLMs with UAVs, with 19.05% fully deploying the technique and 80.95% in the proof-of-concept phase. This indicates that, although the integration of LLMs with UAVs is still in the exploratory stage, some developers have already successfully implemented the initial applications. However, 59.6% of the developers have not yet initiated this project. For those developers who have not implemented the project, we further inquired about the specific reasons for not advancing this. The survey revealed the main challenges.
@@
Challenges Perceived by Developers. The analysis shows that the obstacles developers mention are mainly related to technique, resources, and market factors. In particular, regarding technique, 54.8% of developers believe that existing techniques already meet their needs and therefore are not in a rush to introduce new solutions. This suggests that in some traditional UAV application scenarios, existing techniques are mature enough to complete tasks without the need for more complex LLMs. 51.6% of developers believe that existing large models cannot meet performance requirements, reflecting the current shortcomings of LLMs in terms of performance, computational power, and real-time capabilities, which are still insufficient to meet the high-performance demands in the UAV field. Additionally, 45.2% of developers pointed out the security risks of large models, with some noting that “LLMs are not being used on board in safety-critical applications” and “LLMs should not be on a critical system due to their ‘black box’ nature.” This indicates that data privacy, as well as the transparency and interpretability of models, remain major obstacles to the widespread application of LLMs in the UAV domain.
@@
In addition to the above issues, 25.8% of developers mentioned high costs, noting that the demand for substantial computational resources and specialized hardware reduces cost-effectiveness. 22.6% of developers stated that the lack of training data or fine-tuning methods limits their application, preventing them from effectively utilizing the existing large models. Furthermore, 29.0% of developers believe that a lack of relevant knowledge and technique is another reason this project has not advanced.
[…]
Table 2: Developers’ perceptions on task maturity
@@
Application Maturity Perceived by Developers. To assess developers’ perceptions of the current acceptance and technical maturity of large models in UAV tasks, we invited developers to rate the technical maturity of various tasks of LLMs in UAVs. We used the Likert scale [joshi2015likert] with a scale from 1 to 5, with 1 indicating “completely immature,” 3 representing neutral, and 5 indicating “very mature”. The results, as shown in  2, indicate that the average score for the technical maturity of all tasks is below 3, indicating that while LLMs hold significant application potential in these areas, they also face numerous challenges. Among these, developers considered LLMs to have relatively higher technical maturity in Natural Language Command Parsing and Information Summarization and Report Generation, with scores of 2.89 and 2.96, respectively. These tasks involve basic processing of language and data, and existing LLMs can provide effective support in these areas, resulting in a relatively higher technical maturity. However, for more complex tasks, such as Multi-UAV Task Reasoning and Planning (score 2.40) and Control Logic and Command Generation (score 2.51), the technical maturity was lower. This suggests that, while multi-UAV coordination and autonomous control theoretically have significant potential, there are still many technical obstacles in effectively applying LLMs to these complex tasks in real-world UAV projects. These challenges need further improvement in areas such as accuracy, real-time performance, and collaborative computing capabilities.
@@
Additionally, two developers specifically highlighted potential application scenarios for LLMs in the UAV field, including: supporting users and new developers in setting up UAV systems (e.g., helping configure ArduPilot Wiki) and applications in areas such as natural disaster response, maritime accidents, and search and rescue (SAR), environmental monitoring and hazard detection, and urban safety and public monitoring. This highlights the potential of LLMs in real-world UAV projects, particularly in high-complexity and multi-task processing scenarios, which remain to be further explored.
[…]
@@
Common LLM-UAV Integration Approaches. To gain a deeper understanding of the technical details of integrating LLMs with UAVs, we explored the main integration methods of LLMs in the UAV field. Specifically, we asked developers to select from various integration methods. The survey results show that Hybrid-Onboard Assistance is the most common method, with 82.7% of developers choosing this option. In this approach, LLMs typically do not directly control the UAV’s flight but serve as a supporting tool for task planning, path reasoning, and decision support. This suggests that, currently, LLMs in the UAV field are more focused on enhancing decision-making capabilities to improve the intelligence of task execution, rather than directly intervening in the flight control system.
@@
In comparison, External-Offboard Invocation (i.e., deploying large models on cloud or ground devices to provide decision support via external interfaces) and Internal-Onboard Control (i.e., embedding large models into the UAV’s control system to generate flight control commands directly) were selected by 67.3% and 26.9% of developers, respectively. The application of these two methods is relatively limited and, in practical use, is not as common as the Hybrid-Onboard Assistance approach. This reflects that the application of LLMs in the UAV field is still in the early stages, with the developers more reliant on enhancing decision support rather than direct control.